OpenStudy (anonymous):

Find the area between the graph of y=x^2-2 and the x-axis, between x=0 and x=6.

OpenStudy (luigi0210):

\[\int\limits_{0}^{6}x^2-2\]

OpenStudy (anonymous):

I need to find the area above and below the x axis though and i'm not sure how to do that

OpenStudy (luigi0210):

Do you know what the graph looks like?

OpenStudy (anonymous):

it's a parabola

OpenStudy (anonymous):

 \[\int\limits_{0}^{\sqrt{2}}(x^2-2)dx+\int\limits_{\sqrt{2}}^{6}(x^2-2)dx\]